Solve the following equations:
7m + (19/2) = 13........

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

By transposing (19/2) from LHS to RHS it becomes -19/2
Then,
= 7m = 13–(19/2)
= 7m = (26 - 19)/2
= 7m = 7/2
Now,
Divide both side by 7,
= 7m/7 = (7/2)/7
= m = (7/2) × (1/7)
= m= 1⁄2

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows:
7m + 19/2 = 13
7m = 13 - 19/2 
7m = 7/2
m = 7/14 
m = 1/2
 
Hope it helps!
Thank you ✌️